# Optimism Without Regularization: Constant Regret in Zero-Sum Games

### v1

@@ 1 Introduction
Despite the fact that regularization is essential for no-regret online learning in general adversarial settings, unregularized algorithms are still able to obtain sublinear regret in the case of two-player zero-sum games. Fictitious Play (FP), dating back to Brown, (1951), is the canonical example of such an unregularized algorithm, and it results from both players independently running Follow-the-Leader (FTL).111 FTL is a particular instance of Follow-the-Regularized-Leader (FTRL) with unbounded stepsize $\eta\to\infty$. In the worst case, FTL can have $\Omega(T)$ regret due to its sensitivity to oscillations in adversarially-chosen reward sequences (Shalev-Shwartz et al.,, 2012). However, in zero-sum game settings, the classic result of Robinson, (1951) proved that, under Fictitious Play, the sum of the players’ regrets (henceforth referred to as regret) is indeed sublinear (thus implying time-average convergence to a Nash equilibrium), albeit at the very slow $O(T^{1-1/n})$ rate for $n\times n$ games (Daskalakis and Pan,, 2014).
@@
On the other hand, the past decade has seen regularized learning algorithms exhibit a remarkable success in providing even better $o(\sqrt{T})$ regret guarantees for learning in games. In zero-sum games, optimistic variants of FTRL (including Optimistic Multiplicative Weights and Optimistic Gradient Descent) obtain only constant regret (with respect to the time horizon $T$), implying optimal $O(1/T)$ time-average convergence to Nash (Rakhlin and Sridharan,, 2013; Syrgkanis et al.,, 2015). While such guarantees can be obtained using absolute constant stepsizes (with no dependence on $T$), standard proof techniques (e.g., the RVU bound approach of Syrgkanis et al., (2015)) still crucially require a finite upper bound on stepsize, corresponding to constant magnitudes of regularization. This raises the following, fundamental question:
Is $O(1)$ regret attainable in zero-sum games without regularization (equivalently,
with unbounded stepsizes)? Can variants of Fictitious Play achieve $O(1)$ regret?
@@ 1.1 Our Contributions
In this work, we establish an affirmative answer to the question above. Our main result establishes that, in the case of 2$\times$2 zero-sum games, Optimistic Fictitious Play obtains constant regret:
@@
Optimistic Fictitious Play, using any tiebreaking rule, obtains $O(1)$ regret in all 2$\times$2 zero-sum games with a unique, interior Nash equilibrium.
@@
On the 2$\times$2 Matching Pennies game, Alternating Fictitious Play, using any tiebreaking rule and for nearly all initializations, has regret at least $\Omega(\sqrt{T})$.
Together, our results separate the regret guarantees of using optimism and alternation in the regime of no regularization: while optimism without regularization can obtain optimal $O(1)$ regret (Theorem 3.1), alternation alone is in general insufficient for improving beyond $O(\sqrt{T})$ (Theorem 3.2), the same rate achievable by standard (non-alternating) FP in the 2$\times$2 setting. To this latter point, note that the lower bound of Theorem 3.2 comes in contrast to the improved $O(T^{1/5})$ rate obtainable by Alternating FTRL under a sufficiently small stepsize (Katona et al.,, 2024).
Table 1 summarizes our results and the landscape of regret guarantees in the 2$\times$2 setting for FTRL and FP variants, and Figure 1 shows an example of the empirical regret guarantees of standard, Optimistic, and Alternating FP variants in several games (additional results are presented in Sections 5 and E).
@@
$\eta$ bounded (FTRL) $O(\sqrt{T})$dagger $O(1)$ $O(T^{1/5})$∧
$\eta\to\infty$ (FP) $O(\sqrt{T})$ $O(1)$ $\Omega(\sqrt{T})$⋆
Table 1: Regret guarantees for FTRL and Fictitious Play variants in 2$\times$2 zero-sum games, with our contributions shaded in gray. $\dagger$: Using the standard setting of $\eta=1/\sqrt{T}$ (Shalev-Shwartz et al.,, 2012). : Via the RVU bounds of Syrgkanis et al., (2015). ∧∧: (Katona et al.,, 2024), extending on the prior $O(T^{1/3})$ bound of Wibisono et al., (2022). $\ddagger$: Implicit in the proof of Robinson, (1951). : Theorem 3.1. ⋆⋆: Theorem 3.2.
Figure 1: Empirical regret of standard (FP), Optimistic (OFP), and Alternating (AFP) Fictitious Play in Matching Pennies (from $x^{0}_{1}=(1/3,2/3)$, $x^{0}_{2}=(2/3,1/3)$), on the 15$\times$15 identity matrix (from $x^{0}_{1}=e_{1}$, $x^{0}_{2}=e_{n}$), and on 15$\times$15 generalized Rock-Paper-Scissors (from $x^{0}_{1}=e_{1}$, $x^{0}_{2}=e_{n}$). Each algorithm was run for $T=10000$ iterations using a lexicographical tiebreaking rule. Each subfigure demonstrates the constant empirical regret of OFP compared to the roughly $\sqrt{T}$ regret growth of standard FP and AFP. More experimental details and results are given in Section 5 and Section E.
@@ 1.2 Other Related Work
Optimistic learning in games.
Our work relates to a line of research on the convergence properties of optimistic and extragradient-type algorithms in both normal-form and extensive-form games (Daskalakis et al.,, 2021; Fasoulakis et al.,, 2022; Anagnostides et al., 2022a, ; Zhang et al.,, 2024; Farina et al.,, 2022; Hsieh et al.,, 2022; Anagnostides et al., 2022b, ; Piliouras et al.,, 2022; Anagnostides et al., 2022c, ). As previously noted, a key difference is that these approaches typically rely on constant or time-decreasing step sizes, corresponding to some level of regularization. Recent works have also investigated last-iterate convergence properties of optimistic and extragradient methods (Daskalakis and Panageas,, 2019; Cai et al.,, 2022; Abernethy et al., 2021b, ; Hsieh et al.,, 2020), as well as for variants of regret-matching, including under alternation (Cai et al., 2025a, ). Other works have studied accelerated rates using optimism without regularization in certain Frank-Wolfe-type, convex-concave saddle-point problems (Wang and Abernethy,, 2018; Abernethy et al.,, 2018).
Learning in 2$\times$2 games.
Our work adds to a growing recent literature studying online learning algorithms in 2$\times$2 games: in two-strategy zero-sum games, Bailey and Piliouras, (2019) proved that Online Gradient Descent obtains $O(\sqrt{T})$ regret even with large constant stepsizes. More recent works of Cai et al., (2024) and Cai et al., 2025b establish lower-bounds on the last-iterate and random-iterate convergence rates of Optimistic MWU using a hard 2$\times$2 construction, as well as an $O(T^{1/6})$ upper bound on best-iterate convergence for 2$\times$2 zero-sum games. Chen and Peng, (2020) similarly used a 2$\times$2 construction to establish a general $\Omega(\sqrt{T})$ lower bound on the regret of standard Multiplicative Weights. In a family of two-strategy congestion games, Chotibut et al., (2021) also showed that the iterates of Multiplicative Weights can exhibit formally chaotic behavior.
Fictitious Play.
We refer the reader to the recent results of Daskalakis and Pan, (2014), Abernethy et al., 2021a , and Lazarsfeld et al., (2025) (and the references therein) for background on standard Fictitious Play. Abernethy et al., 2021a also briefly introduced the optimistic variant of Fictitious Play, and they informally conjecture the algorithm to have constant regret on diagonal payoff matrices. The convergence behavior of FP has also been studied in potential games (Monderer and Shapley,, 1996; Panageas et al.,, 2023), near-potential games (Candogan et al.,, 2013), Markov games (Sayin et al.,, 2022; Baudin and Laraki,, 2022), and extensive-form games (Heinrich et al.,, 2015).
@@ 2 Preliminaries
Let $[n]=\{1,\dots,n\}$, let $\Delta_{n}$ be the probability simplex in $\mathbb{R}^{n}$, and let $\{e_{i}\}_{n}=\{e_{i}:i\in[n]\}\subset\Delta_{n}$ denote the set of standard basis vectors in $\mathbb{R}^{n}$, which correspond to vertices of $\Delta_{n}$. For $x\in\Delta_{n}$, we say $x$ is interior if $x_{i}>0$ for all $i\in[n]$.
@@ 2.1 Online Learning in Zero-Sum Games
Let $A\in\mathbb{R}^{m\times n}$ be the payoff matrix for a two-player zero-sum game, and let $T$ be a fixed time horizon. At round $t$, Players 1 and 2 simultaneously choose mixed strategies $x^{t}_{1}\in\Delta_{m}$ and $x^{t}_{2}\in\Delta_{n}$, obtain payoffs $\langle x^{t}_{1},Ax^{t}_{2}\rangle$ and $-\langle x^{t}_{2},A^{\top}x^{t}_{1}\rangle$, and observe feedback $Ax^{t}_{2}$ and $-A^{\top}x^{t}_{1}$, respectively.
@@ 2.2 Fictitious Play and Optimistic Fictitious Play
$x^{t+1}_{1}$ $\;:=\;\mathop{\textnormal{argmax}}_{x\in\{e_{i}\}_{m}}\big{\langle}x,\sum\nolimits_{k=0}^{t}Ax^{k}_{2}+\alpha Ax^{t}_{2}\big{\rangle}$ ($\alpha$-OFP)
$x^{t+1}_{2}$ $\;:=\;\mathop{\textnormal{argmax}}_{x\in\{e_{i}\}_{n}}\big{\langle}x,\sum\nolimits_{k=0}^{t}-A^{\top}x^{k}_{1}-\alpha A^{\top}x^{t}_{1}\big{\rangle}\;.$
@@
[…]
@@ 3 Regret Bounds for Optimistic and Alternating Fictitious Play
As a consequence of the techniques we develop for proving Theorem 3.1, we also establish a lower bound on the regret of Alternating Fictitious Play. In particular, for the corresponding alternating regret $\textnormal{Reg}^{\textnormal{alt}}(T)$ (see Definition D.2), we prove on the 2$\times$2 Matching Pennies game the following:
@@
[…]
@@
[…]
@@
Naively, one might in general hope the invariant $\partial\Psi(y^{t+1})=\partial\Psi(\widetilde{y}^{t+1})$ holds at every timestep under Optimistic FP. However, simple experiments reveal that this is not true: in general $\Psi$ can increase during one step of the algorithm. Thus, understanding when and why such an invariant does hold is still a challenging task that may require leveraging structural properties of the payoff matrix. In the proof of Theorem 3.5, we leverage such properties of 2$\times$2 games to establish sufficient conditions for when the above invariant holds, and this subsequently leads to a constant upper bound on energy.
@@ 4 Bounded Energy Under Optimistic Fictitious Play
As proven by Bailey and Piliouras, (2019), who studied online gradient descent in 2$\times$2 games, for any Fictitious Play or FTRL variant, Assumption 1 holds without loss of generality:
@@
Let $A\in\mathbb{R}^{2\times 2}$ have a unique, interior NE, and let $\{x^{t}\}$ be the iterates of (OFP) on $A$. Then there exists $\widetilde{A}\in\mathbb{R}^{2\times 2}$ satisfying Assumption 1 such that (1) $\widetilde{A}$ and $A$ have the same NE and (2) the iterates $\{\widetilde{x}^{t}\}$ of running (OFP) on $\widetilde{A}$ are identical to $\{x^{t}\}$.
@@
Let $A$ satisfy Assumption 1, and let $\{y^{t}_{1}\}$ and $\{y^{t}_{2}\}$ be the dual payoff vectors of (OFP). Then for every $t\geq 1$, it holds that $y^{t}_{12}=-\rho_{1}\cdot y^{t}_{11}$ and $y^{t}_{22}=-\rho_{2}\cdot y^{t}_{21}$, where $\rho_{1}:=(d-c)/(a-b)>0$ and $\rho_{2}=(d-b)/(a-c)>0$.
The proof of Proposition 4.2 is given in Section C. Importantly, as $\rho_{1},\rho_{2}>0$, observe that $y^{t}_{11}>0\iff y^{t}_{11}>y^{t}_{12}$, and $y^{t}_{21}>0\iff y^{t}_{21}>y^{t}_{22}$ for all times $t\geq 1$. Thus, in the 2$\times$2 setting, the coordinates $y^{t}_{11}$ and $y^{t}_{21}$ encode all information needed to analyze the iterates of Optimistic FP in (3).

### v2

@@ 1 Introduction
Despite the fact that regularization is essential for no-regret online learning in general adversarial settings, unregularized algorithms are still able to obtain sublinear regret in the case of two-player zero-sum games. Fictitious Play (FP), dating back to Brown, (1951), is the canonical example of such an unregularized algorithm, and it results from both players independently running Follow-the-Leader (FTL). [^1] FTL is a particular instance of Follow-the-Regularized-Leader (FTRL) with unbounded stepsize $\eta\to\infty$. In the worst case, FTL can have $\Omega(T)$ regret due to its sensitivity to oscillations in adversarially-chosen reward sequences (Shalev-Shwartz et al.,, 2012). However, in zero-sum game settings, the classic result of Robinson, (1951) proved that, under Fictitious Play, the sum of the players’ regrets (henceforth referred to as regret) is indeed sublinear (thus implying time-average convergence to a Nash equilibrium), albeit at the very slow $O(T^{1-1/n})$ rate for $n\times n$ games (Daskalakis and Pan,, 2014).
@@
On the other hand, the past decade has seen regularized learning algorithms exhibit a remarkable success in providing even better $o(\sqrt{T})$ regret guarantees for learning in games. In zero-sum games, optimistic variants of FTRL (including Optimistic Multiplicative Weights and Optimistic Gradient Descent) obtain only constant regret (with respect to the time horizon $T$), implying optimal $O(1/T)$ time-average convergence to Nash (Rakhlin and Sridharan,, 2013; Syrgkanis et al.,, 2015). While such guarantees can be obtained using absolute constant stepsizes (with no dependence on $T$), standard proof techniques (e.g., the RVU bound approach of Syrgkanis et al., (2015)) still crucially require a finite upper bound on stepsize, corresponding to constant magnitudes of regularization. This raises the following, fundamental question: Is $O(1)$ regret attainable in zero-sum games without regularization (equivalently, with unbounded stepsizes)? Can variants of Fictitious Play achieve $O(1)$ regret?
@@ 1.1 Our Contributions
In this work, we establish an affirmative answer to the question above. Our main result establishes that, in the case of $\times2$ zero-sum games, Optimistic Fictitious Play obtains constant regret:
@@
Optimistic Fictitious Play, using any tiebreaking rule, obtains $O(1)$ regret in all $\times2$ zero-sum games with a unique, interior Nash equilibrium.
@@
On the $\times2$ Matching Pennies game, Alternating Fictitious Play, using any tiebreaking rule and for nearly all initializations, has regret at least $\Omega(\sqrt{T})$.
Together, our results separate the regret guarantees of using optimism and alternation in the regime of no regularization: while optimism without regularization can obtain optimal $O(1)$ regret (Theorem 3.1), alternation alone is in general insufficient for improving beyond $O(\sqrt{T})$ (Theorem 3.2), the same rate achievable by standard (non-alternating) FP in the $\times2$ setting. To this latter point, note that the lower bound of Theorem 3.2 comes in contrast to the improved $O(T^{1/5})$ rate obtainable by Alternating FTRL under a sufficiently small stepsize (Katona et al.,, 2024).
Table 1 summarizes our results and the landscape of regret guarantees in the $\times2$ setting for FTRL and FP variants, and Figure 1 shows an example of the empirical regret guarantees of standard, Optimistic, and Alternating FP variants in several games (additional results are presented in Sections 5 and E).
@@
$\eta$ bounded (FTRL) $O(\sqrt{T})$ $\dagger$ $O(1)$ $O(T^{1/5})$ ∧∧
$\eta\to\infty$ (FP) $O(\sqrt{T})$ $O(1)$ $\Omega(\sqrt{T})$ ⋆⋆
Table 1: Regret guarantees for FTRL and Fictitious Play variants in $\times2$ zero-sum games, with our contributions shaded in gray. $\dagger$: Using the standard setting of $\eta=1/\sqrt{T}$ (Shalev-Shwartz et al.,, 2012). : Via the RVU bounds of Syrgkanis et al., (2015). ∧∧: (Katona et al.,, 2024), extending on the prior $O(T^{1/3})$ bound of Wibisono et al., (2022). $\ddagger$: Implicit in the proof of Robinson, (1951). : Theorem 3.1. ⋆⋆: Theorem 3.2.
Figure 1: Empirical regret of standard (FP), Optimistic (OFP), and Alternating (AFP) Fictitious Play in Matching Pennies (from $x^{0}_{1}=(1/3,2/3)$, $x^{0}_{2}=(2/3,1/3)$), on the $\times15$ identity matrix (from $x^{0}_{1}=e_{1}$, $x^{0}_{2}=e_{n}$), and on $\times15$ generalized Rock-Paper-Scissors (from $x^{0}_{1}=e_{1}$, $x^{0}_{2}=e_{n}$). Each algorithm was run for $T=10000$ iterations using a lexicographical tiebreaking rule. Each subfigure demonstrates the constant empirical regret of OFP compared to the roughly $\sqrt{T}$ regret growth of standard FP and AFP. More experimental details and results are given in Section 5 and Section E.
@@ 1.2 Other Related Work
Optimistic learning in games. Our work relates to a line of research on the convergence properties of optimistic and extragradient-type algorithms in both normal-form and extensive-form games (Daskalakis et al.,, 2021; Fasoulakis et al.,, 2022; Anagnostides et al., 2022a, ; Zhang et al.,, 2024; Farina et al.,, 2022; Hsieh et al.,, 2022; Anagnostides et al., 2022b, ; Piliouras et al.,, 2022; Anagnostides et al., 2022c, ). As previously noted, a key difference is that these approaches typically rely on constant or time-decreasing step sizes, corresponding to some level of regularization. Recent works have also investigated last-iterate convergence properties of optimistic and extragradient methods (Daskalakis and Panageas,, 2019; Cai et al.,, 2022; Abernethy et al., 2021b, ; Hsieh et al.,, 2020), as well as for variants of regret-matching, including under alternation (Cai et al., 2025a, ). Other works have studied accelerated rates using optimism without regularization in certain Frank-Wolfe-type, convex-concave saddle-point problems (Wang and Abernethy,, 2018; Abernethy et al.,, 2018).
Learning in $\times2$ games. Our work adds to a growing recent literature studying online learning algorithms in $\times2$ games: in two-strategy zero-sum games, Bailey and Piliouras, (2019) proved that Online Gradient Descent obtains $O(\sqrt{T})$ regret even with large constant stepsizes. More recent works of Cai et al., (2024) and Cai et al., 2025b establish lower-bounds on the last-iterate and random-iterate convergence rates of Optimistic MWU using a hard $\times2$ construction, as well as an $O(T^{1/6})$ upper bound on best-iterate convergence for $\times2$ zero-sum games. Chen and Peng, (2020) similarly used a $\times2$ construction to establish a general $\Omega(\sqrt{T})$ lower bound on the regret of standard Multiplicative Weights. In a family of two-strategy congestion games, Chotibut et al., (2021) also showed that the iterates of Multiplicative Weights can exhibit formally chaotic behavior.
Fictitious Play. We refer the reader to the recent results of Daskalakis and Pan, (2014), Abernethy et al., 2021a , and Lazarsfeld et al., (2025) (and the references therein) for background on standard Fictitious Play. Abernethy et al., 2021a also briefly introduced the optimistic variant of Fictitious Play, and they informally conjecture the algorithm to have constant regret on diagonal payoff matrices. The convergence behavior of FP has also been studied in potential games (Monderer and Shapley,, 1996; Panageas et al.,, 2023), near-potential games (Candogan et al.,, 2013), Markov games (Sayin et al.,, 2022; Baudin and Laraki,, 2022), and extensive-form games (Heinrich et al.,, 2015).
@@ 2 Preliminaries
Let $n]=\{1,\dots,n\}$, let $\Delta_{n}$ be the probability simplex in $\mathbb{R}^{n}$, and let $\{e_{i}\}_{n}=\{e_{i}:i\in[n]\}\subset\Delta_{n}$ denote the set of standard basis vectors in $\mathbb{R}^{n}$, which correspond to vertices of $\Delta_{n}$. For $x\in\Delta_{n}$, we say $x$ is interior if $x_{i}>0$ for all $i\in[n]$.
@@ 2.1 Online Learning in Zero-Sum Games
Let $A\in\mathbb{R}^{m\times n}$ be the payoff matrix for a two-player zero-sum game, and let $T$ be a fixed time horizon. At round $t$, Players 1 and 2 simultaneously choose mixed strategies $x^{t}_{1}\in\Delta_{m}$ and $x^{t}_{2}\in\Delta_{n}$, obtain payoffs $\langle x^{t}_{1},Ax^{t}_{2}\rangle$ and $\langle x^{t}_{2},A^{\top}x^{t}_{1}\rangle$, and observe feedback $Ax^{t}_{2}$ and $A^{\top}x^{t}_{1}$, respectively.
@@ 2.2 Fictitious Play and Optimistic Fictitious Play
$x^{t+1}_{1}$ $\;=\;\mathop{\textnormal{argmax}}_{x\in\{e_{i}\}_{m}}\big\langle x,\sum\nolimits_{k=0}^{t}Ax^{k}_{2}+\alpha Ax^{t}_{2}\big\rangle$ ($\alpha$-OFP)
$x^{t+1}_{2}$ $\;=\;\mathop{\textnormal{argmax}}_{x\in\{e_{i}\}_{n}}\big\langle x,\sum\nolimits_{k=0}^{t}-A^{\top}x^{k}_{1}-\alpha A^{\top}x^{t}_{1}\big\rangle\;.$
@@
Observe that the $\mathop{\textnormal{argmax}}$ sets in ($\alpha$-OFP) may contain multiple vertices. For this, we assume that the $\mathop{\textnormal{argmax}}$ operator encodes a tiebreaking rule that returns a distinct element. Throughout, we make no assumptions on the nature of the tiebreaking rule, and in general ties can be broken deterministically, randomly, or adaptively/adversarially.
@@ 3 Regret Bounds for Optimistic and Alternating Fictitious Play
As a consequence of the techniques we develop for proving Theorem 3.1, we also establish a lower bound on the regret of Alternating Fictitious Play. In particular, for the corresponding alternating regret $\textnormal{Reg}^{\textnormal{alt}}(T)$ (see Definition D.2), we prove on the $\times2$ Matching Pennies game the following:
@@
$\textbf{For FP:}\quad\Delta\Psi(y^{t})\;\geq\;\big\langle\partial\Psi(y^{t}),J\partial\Psi(\widetilde{y}^{t+1})\big\rangle=\big\langle\partial\Psi(y^{t}),J\partial\Psi(y^{t})\big\rangle=0\;.$ (4)
@@
$\textbf{For Optimistic FP:}\quad\Delta\Psi(y^{t})\leq\big\langle\partial\Psi(y^{t+1}),J\partial\Psi(\widetilde{y}^{t+1})\big\rangle=\big\langle\partial\Psi(y^{t+1}),J\partial\Psi(\widetilde{y}^{t+1})\big\rangle.$ (5)
@@
Naively, one might in general hope the invariant $\partial\Psi(y^{t+1})=\partial\Psi(\widetilde{y}^{t+1})$ holds at every timestep under Optimistic FP. However, simple experiments reveal that this is not true: in general $\Psi$ can increase during one step of the algorithm. Thus, understanding when and why such an invariant does hold is still a challenging task that may require leveraging structural properties of the payoff matrix. In the proof of Theorem 3.5, we leverage such properties of $\times2$ games to establish sufficient conditions for when the above invariant holds, and this subsequently leads to a constant upper bound on energy.
@@ 4 Bounded Energy Under Optimistic Fictitious Play
As proven by Bailey and Piliouras, (2019), who studied online gradient descent in $\times2$ games, for any Fictitious Play or FTRL variant, Assumption 1 holds without loss of generality:
@@
Let $A\in\mathbb{R}^{2\times 2}$ have a unique, interior NE, and let $\{x^{t}\}$ be the iterates of (OFP) on $A$. Then there exists $\widetilde{A}\in\mathbb{R}^{2\times 2}$ satisfying Assumption 1 such that (1) $\widetilde{A}$ and AA have the same NE and (2) the iterates $\{\widetilde{x}^{t}\}$ of running (OFP) on $\widetilde{A}$ are identical to $\{x^{t}\}$.
@@
Let AA satisfy Assumption 1, and let $\{y^{t}_{1}\}$ and $\{y^{t}_{2}\}$ be the dual payoff vectors of (OFP). Then for every $t\geq 1$, it holds that $y^{t}_{12}=-\rho_{1}\cdot y^{t}_{11}$ and $y^{t}_{22}=-\rho_{2}\cdot y^{t}_{21}$, where $\rho_{1}:=(d-c)/(a-b)>0$ and $\rho_{2}=(d-b)/(a-c)>0$.
The proof of Proposition 4.2 is given in Section C. Importantly, as $\rho_{1},\rho_{2}>0$, observe that $y^{t}_{11}>0\iff y^{t}_{11}>y^{t}_{12}$, and $y^{t}_{21}>0\iff y^{t}_{21}>y^{t}_{22}$ for all times $t\geq 1$. Thus, in the $\times2$ setting, the coordinates $y^{t}_{11}$ and $y^{t}_{21}$ encode all information needed to analyze the iterates of Optimistic FP in (3).